### v1

# RoboHanger: Learning Generalizable Robotic Hanger Insertion for Diverse Garments

Abstract
For the task of hanging clothes, learning how to insert a hanger into a garment is crucial but has been seldom explored in robotics. In this work, we address the problem of inserting a hanger into various unseen garments that are initially laid out flat on a table. This task is challenging due to its long-horizon nature, the high degrees of freedom of the garments, and the lack of data. To simplify the learning process, we first propose breaking the task into several stages. Then, we formulate each stage as a policy learning problem and propose low-dimensional action parameterization. To overcome the challenge of limited data, we build our own simulator and create 144 synthetic clothing assets to effectively collect high-quality training data. Our approach uses single-view depth images and object masks as input, which mitigates the Sim2Real appearance gap and achieves high generalization capabilities for new garments. Extensive experiments in both simulation and the real world validate our proposed method. By training on various garments in the simulator, our method achieves a 75% success rate with 8 different unseen garments in the real world.
I Introduction
Laundry is a common and routine household chore that is time-consuming. The entire process involves many steps, including collecting clothes, washing, drying, hanging, unfolding, and folding. Therefore, automating laundry tasks with robots is highly appealing. Researchers have explored how to enable robots to fold and unfold clothes from arbitrary initial configurations [1, 2, 3], or to hang rigid or deformable objects on racks or hooks [4, 5, 6, 7]. However, we have identified that one crucial step in the laundry process — inserting a hanger into clothes — has been rarely explored by researchers.
In this work, we investigate how to use a dual-arm robot to insert a hanger into various unseen garments, as illustrated in Figure LABEL:fig:teaser. Our task begins with the target garment laid flat on the table. The goal is to insert the hanger into the garment and hang it up in the final state. End-to-end imitation learning methods face significant challenges in achieving satisfactory results due to (1) task complexity: inserting a hanger into a garment involves intricate, highly coordinated dual-arm actions in a long horizontal manner; (2) data scarcity: collecting large amounts of real-world demonstrations is inefficient and potentially dangerous, while existing simulators lack the necessary physical accuracy and diversity in suitable garments and hangers; and (3) generalizability: garments vary widely in color, shape, size, and initial configurations, making it difficult for models to generalize across different scenarios.
To address the three aforementioned issues, we propose a method that breaks the task into several stages. An overview of the entire process is provided in Figure 1. For different stages, we leverage both learnable and non-learnable modules to take the best of both worlds. Learnable modules offer more flexible and adaptive actions, while non-learnable modules reduce learning difficulty and improve robustness. We design neural networks to learn actions that need to adapt to new clothing shapes and initial configurations, including pressing, lifting, dragging, and rotating, as detailed in Section III-C. For actions that are less critical to the task’s success, we use fixed motions, such as insertion and pulling actions.
We build a simulator and create over 1,000 hanger and 100 garment assets for data collection. Our simulation environment is built upon Taichi [8], which supports large-scale parallel simulations on GPU. We choose depth and the object’s mask as input instead of raw RGB data to mitigate generalization issues. Depth images provide the necessary geometric information for the policies and are naturally unaffected by the varying colors and patterns of different garments. Extensive experiments in the simulation and real world validate our proposed method. Trained on 120 different garments in the simulation, our method can achieve a final success rate of 75% on 8 different unknown garments in the real world without any fine-tuning with real world data.
II Related Work
II-A Deformable Object Manipulation
Deformable object manipulation is a long-standing challenging topic due to the infinite object states and complex dynamics. Some comprehensive surveys can be found in [9, 10, 11]. One type of deformable object manipulation task involves handling one-dimensional objects such as ropes and cables [12, 13, 14, 15, 16, 17]. 2D clothing-like object manipulation tasks include cloth smoothing [18, 19, 20, 1, 21], cloth folding [22, 23, 2], cloth canonicalization [24, 3], bedding manipulation [25, 26] and bag [27, 28, 29] manipulation. Other works study 3D deformable object manipulation, such as solid object [30, 31] manipulation.
From the aspect of the method, early studies often use the model-based method for planning [32, 33, 34]. These kinds of methods often suffer from inaccurate dynamic models. Some methods have also proposed learning the dynamics model or representations of the deformable objects [35, 36, 37, 38, 39]. Recently, more and more research has explored data-driven approaches for learning policies such as imitation learning [40, 41, 42, 43] and reinforcement learning [6, 44, 45, 46]. Some researchers have also used differentiable simulators for deformable object manipulation tasks [47, 48]. In this work, we employ a data-driven approach, utilizing self-supervised trials to train our value network.
II-B Garment Manipulation
Garment manipulation is an important topic in deformable object manipulation. Folding [49, 50, 2] and unfolding [51, 52, 53, 54, 55] are two basic and widely studied directions. Many works also investigated robotic assisted dressing [56, 57, 58, 59]. Hanging garments is an extremely under-explored area. Some related works include: [6] studied using reinforcement learning to hang a fabric on a rack, [7] learned to hang a knit cap on a coat rack, [60] used a rule based method to manipulate the hanger. In this work, we use learning-based method to enable the robot to insert the hanger into the clothes through the neckline.
III Method
We focus on the task of inserting a hanger into a garment. As illustrated in Figure LABEL:fig:teaser, the initial phase involves a hanger and a garment placed separately on the table in front of the robot. The goal is to control a dual-arm mobile robot that inserts the hanger into the garment through the neckline. The final phase involves the robot lifting the hanger with the garment successfully hanging. To focus on the hanger insertion task, we implement simple heuristic policies for the first and final phases, especially, placing the garment faced up to allow its neckline observable. We also omit the control of the robot’s base, as it is beyond the scope of this work. The robot is manually positioned at the table and remains stationary during execution. Each of the robot’s 7-DoF arms is equipped with parallel grippers, and a depth camera is mounted on its head for perception, as shown in Figure III.
Figure 0: Robot system in the real world. The dual-arm robot is equipped with parallel grippers on both arms and uses a head-mounted camera for observation. The robot’s waist can be controlled to lean forward.
Figure 1: System Overview. Our system takes RGB-D observations. First, we segment the hanger, the garment, and its neckline, and detect the three keypoints of the hanger. The depth map, along with the masks, is then fed into our action primitive networks $Q_{1-4}$. Each network predicts a 2D keypoint $P_{1-4}$ in the image, indicating where the action primitives should operate. These keypoints are first back-projected into 3D space and then sent to the whole-body motion controller to execute the long-horizontal hanger insertion task step-by-step.
We will illustrate the overall pipeline in Section III-A and describe each module including: hanger detection in Section III-B, hanger insertion policy in Section III-C, and action primitive networks in Section III-D. The simulation and data collection process is given in Section III-E and the whole body motion controller is introduced in Section III-F.
III-A System Overview
Our method takes RGB-D images as input, as illustrated in Figure 1. We first detect three keypoints of the hanger: a grasping point, and the left and right endpoints. The grasping point indicates where the gripper should grasp the hanger for subsequent motions, while the left and right endpoints mark the locations where the hanger should be inserted into the target garment. Next, we use GroundedSAM [61] to segment the hanger, garment, and its neckline. When the raw depth data is of low quality, we use [62] to obtain a refined depth map. The refined depth map, along with the masks, is then concatenated and fed into action primitive networks to predict two critical action primitives: press-and-lift and drag-and-rotate, which are used to insert the left and right keypoints of the hanger into the garment, respectively.
III-B Garment Segmentation and Hanger Detection
Given a frame of aligned color and depth images, we first use GroundedSAM [61] to segment and obtain the hanger’s mask. Then, using the camera’s intrinsic and extrinsic parameters, we also obtain the hanger’s point cloud as shown in Figure 1. We design a simple network consisting of a PointNet++ backbone and a 3-layer MLP detection head which outputs three keypoints $\{H_{g},H_{l},H_{r}\}$. $H_{g}$ is the grasping point and $H_{l},H_{r}$ denotes the left and right endpoint of the hanger respectively. To train this network, We synthesize 1,000 meshes of hangers with annotated keypoints.
III-C Hanger Insertion Policy
Action primitives [1, 28, 63] are parameterized end-effector motions specifically designed to accomplish robotic tasks. They are effective for learning and robust in execution. Our hanger insertion policy consists of two key action primitives: press-and-lift and drag-and-rotate, each responsible for inserting the left and right ends of the hanger into the garment. Each action primitive is parameterized by two keypoints. Such action primitives are carefully designed to simplify learning and improve the success rate.
Press-and-lift
As illustrated in Figure 2, after successfully grasping the hanger, we first use the left endpoint $H_{l}$ of the hanger to press the neckline at point $P_{1}$. With the garment pressed by the hanger, we then lift the garment at point $P_{2}$ with the other hand without moving the garment. Finally, we insert the hanger into the garment to a fixed distance. By combining the ’press’ and ’lift’ actions, we can overcome the challenge of the gripper grasping both layers of the garment, which may hinder the insertion of the hanger.
Figure 2: Action primitive: press-and-lift. After grasping the hanger with the right hand, we first press the garment at point $P_{1}$ with the hanger, then lift the garment at point $P_{2}$, and finally insert the hanger to a fixed distance.
Drag-and-rotate
In the second action primitive, as illustrated in Figure 3, we first move the left gripper to the position where the right gripper previously released the hanger and re-grasp the hanger to avoid self-collision between the two arms. Next, we use the right gripper to grasp the clothing at point $P_{3}$ and drag it forward to the right. Then, we use the left gripper to rotate the hanger while moving the right end to $P_{4}$ and aligning the angle. Finally, in real-world experiments, we slightly pull the clothing back to arrange it properly.
Figure 3: Action primitive: drag-and-lift. We first switch to the left hand to re-grasp the hanger, avoiding self-collision during subsequent actions. Next, we use the right hand to drag the garment and rotate the hanger into it. Finally, we pull the garment to arrange it properly.
Success criteria and reward
For each action primitive, the success criterion is whether the hanger’s endpoint ($H_{l}$ for the first action primitive and $H_{r}$ for the second) is inserted into the garment. Similarly, if both endpoints are inserted, the task is considered successfully completed. This success criterion also serves as a sparse reward for training our action primitive networks. In simulation, we cast multiple rays randomly from the endpoint in all directions. If the proportion of rays intersecting with the garment’s mesh exceeds a threshold (0.95 for the left endpoint and 0.9 for the right in our experiments), we consider the endpoint to be covered by the garment.
III-D Action Primitive Networks
We use separate value networks to learn the parameters (operation points) for the two action primitives. Specifically, each network learns an action value map using a UNet architecture [64], as illustrated in Figure 1. Each pixel in this map corresponds to the action value of operating on that point. The network input, denoted as $O$, consists of a top-down depth map and object masks. Specifically, we use the garment mask and neckline mask when predicting $P_{1}$ and $P_{2}$ for the press-and-lift action, and the garment mask and hanger mask when predicting $P_{3}$ and $P_{4}$ for the drag-and-rotate action.
To train the four value networks, we simulate over 10,000 trajectories and their corresponding ground truth rewards. Each data entry is of the form $(O,P_{1},P_{2},R)$ where $P_{1}=(X_{1},Y_{1})$ and $P_{2}=(X_{2},Y_{2})$ correspond to each pixel in the output value map, and $R$ is the sparse reward received after executing these actions. The second value network $Q_{2}$ can be directly supervised using the sparse reward $R$ at the end of the execution of both key steps:
$\mathcal{L}_{2}=BCE(Q_{2}[P_{2};O,P_{1}],R)$ (1)
where $BCE$ denotes the binary cross-entropy loss and []delimited-[][$\cdot$][ ⋅ ] is the indexing operation. The input to the $Q_{2}$ is the observation $O$ concatenated with the one-hot map of $P_{1}$. The $Q_{2}$ network serves two purposes here: first, it can predict the value map for $P_{2}$ given $P_{1}$; second, the maximum of its output indicates the action value of point $P_{1}$. Therefore, we propose to use the maximum of the second network’s output $\hat{R}=(\max\{Q_{2}[\cdot;O,P_{1}]\}>0.5)$ as pseudo label for $Q_{1}$. Here $Q_{2}$ is the output after the Sigmoid function. The final loss function for the first network $Q_{1}$ is:
$\mathcal{L}_{1}=BCE(Q_{1}[P_{1};O],\hat{R})$ (2)
We apply a similar data collection and training strategy to train the value networks $Q_{3}$ and $Q_{4}$. When executing these action primitives, we back-project pixels into three-dimensional coordinates using depth value.
III-E Data Collection
We observe that existing simulators, such as PyFlex [65] and MuJoCo [66], either lack the accuracy or efficiency needed for our task. Their main issue is that the self-collisions in the clothing is not well handled. Therefore, we build our simulation environment based on Taichi [8].
Simulation environment and assets
We use FEM and implicit time integration with a single-step Newton iteration [67] for physical simulation. Cloth self-collision and cloth-to-rigid-body collision are handled with a cubic energy function. Cloth self-intersection is prevented using Continuous Collision Detection (CCD). The collision force between cloth and rigid bodies is computed through the Signed Distance Function (SDF) of the latter. We do not perform dynamics simulation or collision handling for rigid bodies and articulated bodies; instead, we compute the object’s position and velocity using kinematic relationships. We render depth images with SAPIEN [68] and design the 3D models of clothes and hangers (without texture) heuristically.
Domain randomization
We randomize the Young’s modulus and density of the clothing, as well as the friction coefficient of the garment in the simulation. Additionally, we randomize the camera’s intrinsic and extrinsic parameters to compensate for calibration errors when the system is deployed in the real world.
Data collection strategy
We use an iterative data collection approach. Specifically, we first design a state-based heuristic policy to collect trajectories, and use this data to train our networks for initialization. The heuristic policy attempts to find an action that meets all of our predefined rules through sampling. These rules are based on geometric information of the current clothing mesh, such as the number of layers of the garment near the grasping point. Next, we use the network policy to collect more data in the simulator, incorporate this online data into the dataset, and iteratively finetune the network.
III-F Whole-body Motion Controller
Since our hanger insertion task requires a large working space — with some actions necessitating the agent to lean forward without colliding with the table — we implement a custom whole-body controller to ensure the action space does not become a bottleneck. Specifically, after specifying the position of the robotic arm’s end effector, we only constrain the grippers’ fingertips to remain parallel to the table, leaving the other two rotational degrees of freedom unrestricted. Additionally, for steps that require an even larger workspace, we adjust the waist joint angles to allow the robot to lean forward. These angles are computed using a linear function based on the target positions of the left and right end effectors.
IV Experiment
IV-A Experiment Setup
Hanger. We find that using parallel grippers makes it difficult to reliably grasp thin hangers. Therefore, we employ custom-designed 3D-printed hangers to facilitate grasping. The primary difference between our hangers and regular ones is that our design includes a section specifically made for secure gripping.
Metrics. We use the success rate of inserting the hanger, $S$, to evaluate the performance of the policy. $S$ can be further divided into: the success rate of inserting the left endpoint, $S_{1}$; and the success rate of inserting the right endpoint after successfully inserting the left endpoint, $S_{2}$. In our policy, the left end is always inserted first, followed by the right end. Therefore, $S=S_{1}\times S_{2}$.
[…]
IV-B Comparison with Baselines in Simulation
We first compare our methods with several other baselines in simulation. Among the 144 synthetic clothing meshes, we randomly divided them into 120 training garments and 24 testing garments. The hangers used for training and testing are the ones shown in Figure 4 (a.1) and (a.2).
Methods Input Output S1(%) $\uparrow$ S2(%) $\uparrow$ S(%) $\uparrow$
Ours V Prim $88.7$ $\mathbf{95.9}$ $\mathbf{85.1}$
Heuristic S Prim $59.1$ $75.4$ $44.5$
IL V Prim $85.8$ $91.7$ $78.7$
SAC [69] V Prim $37.9$ $56.0$ $21.2$
ACT [43] V Joint $\mathbf{94.8}$ $67.0$ $63.5$
DP [70] V Joint $82.8$ $59.7$ $49.5$
TABLE I: Comparison with baseline policies in simulation.
We present the comparisons in Table I. ”Input” refers to the type of input used by the policy, which can be either visual input (V) or state input (S). ”Output” refers to the type of output generated by the policy, which can either be action primitives (Prim) or joint angles (Joint).
”Heuristic” represents a heuristic algorithm that we manually designed based on the current mesh of the clothing.
”IL” stands for imitation learning, which is also based on action primitives and UNets. We treat the task as a multi-classification task (selecting a pixel in the image) using cross-entropy loss and only retain positive samples for training the UNets.
”SAC” stands for Soft Actor-Critic algorithm [69]. In SAC, the input to the value network consists of both the observation and the action. The action is first transformed into an image where each pixel value represents the inverse of the distance to that action. This action image is then concatenated with the observation and fed into a CNN. The action network is also a CNN that directly outputs $(X_{i},Y_{i})$. Compared to our method, SAC requires learning an additional action network, while our method directly finds the maximum position on the output of the value network.
”ACT” and ”DP” (diffusion policy) are methods proposed by [43] and [70]. We directly use data generated by heuristics to train an end-to-end imitation learning policy. The network’s input consists of the raw depth images from the head camera as well as the masks of the clothing and the hanger.
Our method performs best in simulation and achieves an overall 85.1% success rate. ACT performs excellently when inserting the first endpoint, but due to accumulated errors, it performs poorly when inserting the second endpoint. Similar issues are observed in other end-to-end imitation learning baselines such as diffusion policy. We also observe that SAC performs worst probably due to the additional action network not sufficiently trained.
IV-C Ablation Studies in Simulation
We provide the ablation results in Table II to validate our design choices.
S1(%) $\uparrow$ S2(%) $\uparrow$ S(%) $\uparrow$
Ours $\mathbf{88.7}$ $\mathbf{95.9}$ $\mathbf{85.1}$
w/o online data $86.5$ $94.7$ $81.8$
w/o neckline mask $85.6$ $94.2$ $80.6$
TABLE II: Ablation study in simulation.
”w/o online data” indicates training the network using only offline data without online data, under the same data quantity. Fine-tuning with online data leads to an improvement in the final success rate. ”w/o neckline mask” indicates training the policy without using the neckline mask. It is evident that the neckline mask also contributes to an increase in the final success rate.
IV-D Comparison with Baselines in Real World
In real-world experiments, we test 8 different garments and Hanger 1, as shown in Figure 4. For a fair comparison, we use the robot to initialize the garment configuration. Additionally, to evaluate our policy’s performance under different initializations, we also report the policy’s success rate with human initialization.
Output S1(%) $\uparrow$ S2(%) $\uparrow$ S(%) $\uparrow$
Ours Prim $\mathbf{93.8}$ $80.0$ $\mathbf{75.0}$
IL Prim $75.0$ $\mathbf{83.3}$ $62.5$
ACT Joint $12.5$ $0.0$ $0.0$
Fix Prim $75.0$ $41.7$ $31.2$
  Ours(H) Prim $93.8$ $85.3$ $80.0$
TABLE III: Experiments in the real world.
We present the performance of our policy compared to other methods on the real robot in Table III. We include ”IL” (imitation learning), which performed best among baseline policies in simulation; ”ACT” (a policy that directly outputs joint angles); and ”Fix” (a policy outputs hard coded parameters of both action primitives) as baselines. Our method outperforms all baseline methods in terms of the final success rate which demonstrates the efficacy of our method.
Although the ACT policy shows a certain level of success in simulation, with a particularly high success rate when inserting the first endpoint, it suffers from a severe sim-to-real gap and almost never succeeds in the real world.
Fixed policy is a manually tuned policy for the test garments. It can be observed that the fixed policy still shows high success rates for inserting the left endpoint because the initial position of the clothes is relatively fixed in our experiments. However, when inserting the right endpoint, the success rate of the fixed policy drops significantly due to unforeseen deformations of the clothes during the left endpoint insertion.
In the last row of the table (”Ours(H)”), we further test the performance of our policy when the garment is initialized by a human. When the garment is initialized by the robot, the success rate shows only a slight decrease compared to human initialization, demonstrating the robustness of our policy to the initialization process.
IV-E Generalization of Hangers in Real World
To further study the generalizability of our method in terms of different hangers, we report the performance of our policy when using different hangers in Table IV. Here, Hanger 2 is a hanger seen during training but without a crossbar. Hanger 3 and Hanger 1 have the same geometry, but due to difference settings in the 3D printing process, Hanger 3’s surface is much rougher. Hanger 4 is a hanger that is not seen during training.
S1(%) $\uparrow$ S2(%) $\uparrow$ S(%) $\uparrow$
Hanger 1 $\mathbf{93.8}$ $\mathbf{80.0}$ $\mathbf{75.0}$
Hanger 2 $81.2$ $23.1$ $18.8$
Hanger 3 $75.0$ $75.0$ $56.2$
Hanger 4 $\mathbf{93.8}$ $66.7$ $62.5$
TABLE IV: Comparison of performance with different hangers in the real world.
Although Hanger 2 was seen during training, there is a significant performance drop when inserting the right endpoint. As shown in Figures 5(a) and 5(b), these two images compare the state when inserting the right endpoint using different hangers. The presence of a crossbar helps the hanger separate the upper and lower layers of the clothing, making insertion easier. When using Hanger 3, as shown in Figures 5(c) and 5(d), if the hanger has too much friction, it may drag the clothes while moving, causing the clothes to become very wrinkled, which in turn reduces the success rate. When using Hanger 4, the complexity of the hanger’s shape also affects the final success rate.
Figure 5: Results with different hangers in the real world. In the real world, issues may arise when changing hangers. Figure (a) and Figure (c) show results with Hanger 1. As shown in Figure (b), when switching to Hanger 2, it becomes difficult to separate the upper and lower layers of the clothing while inserting the right end. Figure (d) illustrates that with Hanger 3, which has higher friction, the clothing tends to get wrinkled during the insertion of the left end, causing the garment to become stuck in a worse state.
V Limitation
For garments with less elasticity, such as shirts, more precise force feedback control might be required to avoid damage to the hanger and garment. Another challenge could arise when applying our action primitives to irregularly shaped garments, such as a shirt with a hood.
VI Conclusion
In this work, we investigate the task of inserting a hanger into the neckline of clothes. For such a challenging but rarely studied task, we propose a simulation environment and data collection pipeline to address it. We demonstrate the efficiency of training the algorithm in a simulated environment, as well as its effectiveness and generalizability in the real world. Our pipeline facilitates efficient demonstration collecting for future robotic foundation models that require large amounts of diverse data. Future work may involve relaxing the assumptions for this task and extending the method to handle a wider variety of garments.